OpenStudy (anonymous):

Write the equation of the line wich passes through (2,-3) and is perpendicular to y=4x+7 in standard form.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (lncognlto):

The slope of y=4x+7 is 4. Therefore the slope of the perpendicular line = -1/4, because the slopes of perpendicular lines multiplied together equal -1. Substituting m = -1/4 and (2, -3) into y - y1 = m(x - x1): therefore: y -(-3) = -1/4(x - 2) tf 4y = -x + 2 - 3 tf x + 4y = -1